# A Single-Loop Gradient Algorithm for Pessimistic Bilevel Optimization via Smooth Approximation

### v1

@@ C.1 Auxiliary Lemmas
This completes the proof. $\square$
@@
$\|\nabla\phi_{k}(x)-\nabla\phi_{k}(x^{\prime})\|$ $=\|\nabla_{x}\psi_{k}(x,u_{k}^{*}(x))-\nabla_{x}\phi_{k}(x^{\prime},u_{k}^{*}(x^{\prime}))\|$ (48)
@@
where the final inequality follows from Lemma C.2. $\square$
@@
Finally, combining the above three inequalities with (50), we arrive at the desired inequality. $\square$
@@
where the last inequality follows from Lemma C.1. The conclusion follows by combining the above inequality with (55) and (54). $\square$
@@
Furthermore, the summability in (60) implies
$\lim_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|/\alpha_{k}=\lim_{k\rightarrow\infty}\|u^{k}-u_{k}^{*}(x^{k})\|=0.$
@@
$\quad\|\mathcal{P}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k})-\mathcal{P}_{\mathcal{X}}(x_{k}-\alpha_{k}d_{x_{k}})\|$
$\leq\alpha_{k}\|\nabla\phi_{k}(x^{k})-d_{x_{k}})\|$
$=\alpha_{k}\|\nabla_{x}\psi_{k}(x^{k},y_{k}^{*}(x^{k}),z_{k}^{*}(x^{k}))-\nabla_{x}\psi_{k}(x^{k},y^{k+1},z^{k+1})\|$
$\leq\alpha_{k}(L_{F}+2\rho_{k}L_{f})\|u^{k+1}-u_{k}^{*}(x^{k})\|$
$\leq\alpha_{k}(L_{F}+2\rho_{k}L_{f})\|u^{k}-u_{k}^{*}(x^{k})\|,$
where the first inequality follows from the nonexpansiveness of the projection operator $\mathcal{P}_{\mathcal{X}}$, and the last inequality follows from Lemma C.1. Since $\|u^{k}-u_{k}^{*}(x^{k})\|\rightarrow 0$ as $k\rightarrow\infty$, and $\alpha_{k}\rho_{k}=\alpha_{0}\rho_{0}k^{-s+p}\rightarrow 0$ as $k\rightarrow\infty$, it follows that
$\lim_{k\rightarrow\infty}\|\mathcal{P}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k})-\mathcal{P}_{\mathcal{X}}(x_{k}-\alpha_{k}d_{x_{k}})\|=0.$
Furthermore, we have previously shown that:
$\lim_{k\rightarrow\infty}\frac{1}{\alpha_{k}}\|x^{k}-\mathcal{P}_{\mathcal{X}}(x_{k}-\alpha_{k}d_{x_{k}})\|=\lim_{k\rightarrow\infty}\frac{1}{\alpha_{k}}\|x^{k}-x^{k+1}\|=0.$
Combining this with the convergence of the projected gradients yields:
$\lim_{k\rightarrow\infty}\frac{1}{\alpha_{k}}\|x^{k}-\mathcal{P}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k})\|=0.$
@@
$\lim_{k\rightarrow\infty}\frac{1}{\alpha_{k}}\|x^{k}-\mathcal{P}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k})\|=0.$
This condition, together with the Lipschitz continuity of $\nabla\phi_{k}$ established in Lemma C.4, implies that $x^{k}$ is an approximate stationary point for $\phi_{k}$. Standard arguments then show that for any $\epsilon>0$, there exists $K>0$ such that for all $k\geq K$, there is a $\delta_{k}>0$ such that
$\phi_{k}(x)\geq\phi_{k}(x^{k})-\epsilon\|x-x^{k}\|,\qquad\forall x\in\mathbb{B}_{\delta_{k}}(x^{k})\cap X.$
According to Lemma B.5, for each $k$:
$\phi_{k}(x^{k})\geq\phi(x^{k})-\frac{\sigma_{k}}{2}\|y^{*}(x^{k})\|^{2}$ (63)
where $y^{*}(x)=\arg\max_{y\in\mathcal{S}(x)}F(x,y)$. By Lemma B.4, there exists $M_{y^{*}}>0$ such that for any $k$,
$\|y^{*}(x^{k})\|\leq M_{y^{*}}.$
Since $\sigma_{k}\rightarrow 0$ as $k\rightarrow\infty$, we can obtain from (63) that for any $\epsilon>0$, there exists $K>0$ such that for each $k\geq K$,
$\phi_{k}(x^{k})\geq\phi(x^{k})-\epsilon.$
Furthermore, by Lemma C.9, for any $\epsilon>0$, there exists $K>0$ such that for any $k\geq K$,
$\phi_{k}(x)\leq\phi(x)+\epsilon,\qquad\forall x\in X.$
Combining these inequalities yields that for any $\epsilon>0$ and $\tilde{\epsilon}>0$, we can find $K>0$ such that for each $k\geq K$, there exists $\delta_{k}>0$ such that
$\phi(x)\geq\phi(x^{k})-\epsilon\|x-x^{k}\|-\epsilon,\qquad\forall x\in\mathbb{B}_{\delta_{k}}(x^{k})\cap X,$
completing the proof. $\square$

### v4

@@ C.1 Auxiliary Lemmas
This completes the proof.
$\square$
@@
$\|\nabla\phi_{k}(x)-\nabla\phi_{k}(x^{\prime})\|$ $=\|\nabla_{x}\psi_{k}(x,u_{k}^{*}(x))-\nabla_{x}\psi_{k}(x^{\prime},u_{k}^{*}(x^{\prime}))\|$ (48)
@@
where the final inequality follows from Lemma C.2.
$\square$
@@
Finally, combining the above three inequalities with (50), we arrive at the desired inequality.
$\square$
@@
where the last inequality follows from Lemma C.1. The conclusion follows by combining the above inequality with (55) and (54).
$\square$
@@
Furthermore, since $b_{k}\beta_{k}\sigma_{k}/(L_{F}+2\rho_{k}L_{f})=O(1/K^{-7p-7q})$ and 7p7q<17p-7q<1, the summability in (60) implies
$\underset{k\rightarrow\infty}{\lim\inf}\;\max\{\|x^{k+1}-x^{k}\|/\alpha_{k},\;(L_{F}+2\rho_{k}L_{f})\|u^{k}-u_{k}^{*}(x^{k})\|,\|u^{k}-u_{k}^{*}(x^{k})\|\}=0.$
@@
$\quad\frac{1}{\alpha_{k}}\|\mathrm{Proj}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k})-\mathrm{Proj}_{\mathcal{X}}(x_{k}-\alpha_{k}d_{x}^{k})\|$
$\leq\|\nabla\phi_{k}(x^{k})-d_{x}^{k})\|$
$=\|\nabla_{x}\psi_{k}(x^{k},y_{k}^{*}(x^{k}),z_{k}^{*}(x^{k}))-\nabla_{x}\psi_{k}(x^{k},y^{k+1},z^{k+1})\|$
$\leq(L_{F}+2\rho_{k}L_{f})\|u^{k+1}-u_{k}^{*}(x^{k})\|$
$\leq(L_{F}+2\rho_{k}L_{f})\|u^{k}-u_{k}^{*}(x^{k})\|,$
where the first inequality follows from the nonexpansiveness of the projection operator $\mathrm{Proj}_{\mathcal{X}}$, and the last inequality follows from Lemma C.1. Combining this with the previous equality yields:
$\underset{k\rightarrow\infty}{\lim\inf}\;\frac{1}{\alpha_{k}}\|x^{k}-\mathrm{Proj}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k}))\|$
$=\;$ $\underset{k\rightarrow\infty}{\lim\inf}\;\frac{1}{\alpha_{k}}\|x^{k}-\mathrm{Proj}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k}))+\mathrm{Proj}_{\mathcal{X}}(x_{k}-\alpha_{k}d_{x}^{k})-x^{k+1}\|$
$\leq\;$ $\underset{k\rightarrow\infty}{\lim\inf}\;\left(\frac{1}{\alpha_{k}}\|\mathrm{Proj}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k})-\mathrm{Proj}_{\mathcal{X}}(x_{k}-\alpha_{k}d_{x}^{k})\|+\|x^{k}-x^{k+1}\|\right)$
$\leq\;$ $\underset{k\rightarrow\infty}{\lim\inf}\;(L_{F}+2\rho_{k}L_{f})\|u^{k}-u_{k}^{*}(x^{k})\|+\|x^{k}-x^{k+1}\|$
$\leq\;$ $\underset{k\rightarrow\infty}{\lim\inf}\;2\max\{\|x^{k+1}-x^{k}\|/\alpha_{k},\;(L_{F}+2\rho_{k}L_{f})\|u^{k}-u_{k}^{*}(x^{k})\|\}$
$=\;$ $0.$
@@
$\underset{k\rightarrow\infty}{\lim\inf}\;\frac{1}{\alpha_{k}}\|x^{k}-\mathrm{Proj}_{\mathcal{X}}(x_{k}-\alpha_{k}\nabla\phi_{k}(x^{k})\|=0.$
Thus, we can find a subsequence $\{x^{k_{j}}\}$ such that $\lim_{j\to\infty}\|x^{k_{j}}-\mathrm{Proj}_{\mathcal{X}}(x_{k_{j}}-\alpha_{k_{j}}\nabla\phi_{k_{j}}(x^{k_{j}})\|=0.$ This condition, together with the Lipschitz continuity of $\nabla\phi_{k_{j}}$ established in Lemma C.4, implies that $x^{k_{j}}$ is an approximate stationary point for $\phi_{k_{j}}$. Standard arguments then show that for any $\epsilon>0$, there exists $K_{0}>0$ such that for all ${k_{j}}\geq K_{0}$, there is a $\delta_{j}>0$ such that
$\phi_{k_{j}}(x)\geq\phi_{k_{j}}(x^{k_{j}})-\epsilon\|x-x^{k_{j}}\|,\qquad\forall x\in\mathbb{B}_{\delta_{j}}(x^{k_{j}})\cap X.$
According to Lemma B.5, for each $k_{j}$}:
$\phi_{k_{j}}(x^{k_{j}})\geq\phi(x^{k_{j}})-\frac{\sigma_{k_{j}}}{2}\|y^{*}(x^{k_{j}})\|^{2}$ (63)
where $y^{*}(x)=\arg\max_{y\in\mathcal{S}(x)}F(x,y)$. By Lemma B.4, there exists $M_{y^{*}}>0$ such that for any $k_{j}$},
$\|y^{*}(x^{k_{j}})\|\leq M_{y^{*}}.$
Since $\sigma_{k_{j}}\rightarrow 0$ as $k_{j}$$\rightarrow\infty$, we can obtain from (63) that for any $\tilde{\epsilon}>0$, there exists $K_{0}>0$ such that for each ${k_{j}}\geq K$,
$\phi_{k_{j}}(x^{k_{j}})\geq\phi(x^{k_{j}})-\frac{\tilde{\epsilon}}{2}.$
Furthermore, by Lemma C.9, for any $\tilde{\epsilon}>0$, there exists $K_{0}>0$ such that for any ${k_{j}}\geq K_{0}$,
$\phi_{k_{j}}(x)\leq\phi(x)+\frac{\tilde{\epsilon}}{2},\qquad\forall x\in X.$
Combining these inequalities yields that for any $\epsilon>0$ and $\tilde{\epsilon}>0$, we can find $K_{0}>0$ such that for each ${k_{j}}\geq K_{0}$, there exists $\delta_{j}>0$ such that
$\phi(x)\geq\phi(x^{k_{j}})-\epsilon\|x-x^{k_{j}}\|-\tilde{\epsilon},\qquad\forall x\in\mathbb{B}_{\delta_{k_{j}}}(x^{j})\cap X.$
Since for all $K>K_{0}$, we can find some $k_{j}>K$, the proof is completed. $\square$